A solid metal model of a rocket is made by fastening a cone of vertical height 49 cm and base radius 18 cm to a circular cylinder of length 192 cm and radius 18 cm. If the mass of the model is 2145 kg, calculate the density of the metal in kg/m^3, giving your answer correct to the nearest whole number. :)
Answer by mananth(16946)

A solid metal model of a rocket is made by fastening a cone of vertical height 49 cm and base radius 18 cm to a circular cylinder of length 192 cm and radius 18 cm. If the mass of the model is 2145 kg, calculate the density of the metal in kg/m^3, giving your answer correct to the nearest whole number. :)
...
height of cone = 49cm= 0.49m
base radius = 18 cm = 0.18 m
...
Volume of cone V= 1/3 * pi*r^2*h
V= 1/3 * pi*(0.18)^2*0.49
V= 0.0166 m^3
...
Volume of cylinder = pi*r^2*h
r=18cm=0.18m
height = 192 cm = 1.92 m
Volume = pi*(0.18)^2*1.92
V= 0.195 m^3
Total volume = 0.0166+0.195
Volume = 0.211 m^3...
Density = Mass/Volume
Density = 2145/0.211
Density =10,165kg/m^3
...
m.ananth@hotmail.ca